### v2

# FIVB ranking: Misstep in the right direction

Abstract
This work presents and evaluates the ranking algorithm that has been used by Fédération Internationale de Volleyball (FIVB) since 2020. The prominent feature of the FIVB ranking is the use of the probabilistic model, which explicitly calculates the probabilities of the future matches results using the estimated teams’ strengths. Such explicit modeling is new in the context of official sport rankings, especially for multi-level outcomes, and we study the optimality of its parameters using both analytical and numerical methods. We conclude that from the modeling perspective, the current thresholds fit well the data but adding the home-field advantage (HFA)would be beneficial. Regarding the algorithm itself, we explain the rationale behind the approximations currently used and show a simple method to find new parameters (numerical score) which improve the performance. We also show that the weighting of the match results is counterproductive.
1 Introduction
The ranking of teams/players is one of the fundamental problems in competitive sports. It is used, for example, to declare the champion or to promote and relegate teams between leagues, and, in international competitions, to establish the composition of groups e.g., in the qualification rounds of the FIFA World Cup. In other words, the ranking is a tool that allows the governing bodies to manage competitions by “fairly” evaluating the teams. In general terms, the ranking is meant to reflect the relative strength of the teams/players, and may be used for a quick assesment of the competitive landscape and, more consequentially, for tournament design where the strongest teams are scheduled to play in different group phase (Csató, 2024). The formal evaluation of the ranking is based on its ability to predict the results of future matches (Csató, 2024).
In this work, we present and evaluate the ranking algorithm, used by Fédération Internationale de Volleyball (FIVB). Our study is motivated by the modern approach adopted by FIVB, where six-level outcomes of the volleyball matches have explicit probabilistic models defined by FIVB (2024). To our knowledge, among main international sports, [^1] Such as football, cricket, field hockey, tennis, volleybal, table tennis, golf, basketball, ice hockey, or tennis. this is the first officially adopted ranking algorithm to use such an approach and, merely due to this fact, deserves attention in sport analytics. Of course, international volleyball is also a popular sport, and understanding its ranking strategy is interesting on its own merit.
The FIVB ranking adopted in 2020 can be classified as power-ranking, where teams are assigned a real-valued parameter called skills (also known as strength or power) and the teams are ranked (ordered) by sorting the skills. This approach departs from the more conventional ranking based on counting of the points associated with the results of the matches, and is often considered to be more “fair”.
The power-ranking approach was already adopted by Fédération Internationale de Football Association (FIFA) to rank the Men’s and Women’s teams. The analysis of Szczecinski and Roatis (2022) revealed its weaknesses, where one of the criticisms was that the FIFA ranking, being based on the Elo ranking (Elo, 2008), inherits its main drawback, i.e., the lack of an explicit probabilistic model of the match’s outcomes (Szczecinski and Djebbi, 2020). From a statistical perspective, this lack of forecasting capability is, indeed, a significant drawback which makes it difficult to evaluate the ranking objectively. In this regard, the FIVB ranking proposes a radical and modern approach: each of the six outcomes of the volleyball match is assigned a probability that is calculated from the skills (known before the match) using the Cumulative Link (CL) model (Tutz, 2012, Ch. 9.1).
The objective of this work is to reveal the assumptions and simplifications used to derive the FIVB algorithm and to evaluate how well they are applied. In particular, we want to: i) propose the evaluation methodology of the FIVB ranking by casting it in a statistical framework, ii) show approximations used to derive the algorithm, and to iii) assess the optimality of the algorithm’s parameters. In other words, we want to explain how and why the algorithm is working and whether it can be improved in the current general formulation. In this regard, our work follows the approach of Szczecinski and Roatis (2022) which focused on evaluation and understanding of the currently used FIFA ranking. That is, we do not want to propose/analyze new models or algorithms which would be rather in the spirit of many previous works in the area of sport ranking, e.g., in (Karlis and Ntzoufras, 2008), (Egidi et al., 2018), (Ntzoufras et al., 2019), (Gabrio, 2020), (Lasek and Gagolewski, 2021), (Szczecinski, 2022), or (Macrì Demartino et al., 2024).
Since there are many sports with multi-level outcomes (i.e., taking more than two possible values), providing an understanding of how ranking algorithms may be constructed in such cases will be useful beyond the context of the FIVB ranking. We note that the previous works, e.g., (Egidi and Ntzoufras, 2019) or (Ntzoufras et al., 2019), already addressed the issue of modeling in volleyball. In our work, however, we analyze the ranking algorithm currently used by FIVB which, as we show, is not straightforwardly deduced from the model.
This work is organized as follows: in Sec. 2, we cast the ranking in the inference context, where the goal is to estimate the skills from the outcomes of the matches. This allows us to understand the approximate relationship between the probabilistic model and the FIVB ranking algorithm as such. The parameters of the model are assessed in Sec. 3, where, both the analytical and numerical approaches are applied to evaluate the importance of the model thresholds, the numerical scores used in the FIVB algorithms, the role of the home-field advantage (HFA), and the utility of weights associated with matches’ categories. The parameters obtained in Sec. 3, are used in Sec. 4 to assess the performance of the real-time ranking using the FIVB international matches.
We terminate the work in Sec. 5 summarizing the findings and showing the recommended changes to the FIVB ranking algorithm. Overall, we conclude that, from the statistical perspective, using an explicit probabilistic model to build a ranking algorithm is a step in the right direction. On the other hand, we qualify it as misstep because many approximations lead to sub-optimality, which might have been easily avoided.
The repository https://github.com/brbalab/FIVB contains code/data from which the results can be reproduced.
2 Model and ranking
Consider the scenario, where, of the $M$ teams, two are selected to face each other in a match. The matches are indexed with $t=1,2,\ldots,T$. Teams are indexed with a pair $(i_{t},j_{t})$, where $i_{t},j_{t}\in\{1,\ldots,M\}$. The team $i_{t}$ is called the home-team, and the team $j_{t}$ the away-team. We keep this naming convention even when a match is played on a neutral venue. [^2] A venue is neutral when, during an international tournament, the match is played in the country which is not the one of the team $i_{t}$ or $j_{t}$.
The outcome of the $t$-th match $y_{t}\in\mathcal{Y}$ is ordinal in nature, with meaning such as “importance”, which has no numerical value but may be ordered, and, for convenience, we index it with natural numbers $\mathcal{Y}=\{0,\ldots,L-1\}$ in decreasing order, i.e., the outcome $y=0$ is the most important and $y=L-1$ is the least important one, where the importance is evaluated from the point of view of the home team $i_{t}$. Quite naturally, the most important outcome for the home team is the least important for the away team.
In particular, volleyball matches, which we will be interested in, produce $L=6$ possible outcomes “3-0”, “3-1”, “3-2”, “2-3”, “1-3”, “0-3”, where “$k$-$l$” means that the home team won $k$ sets and the away team won $l$ sets; i.e., the first three results mean that the home team won. For the purpose of the derivations, we use indices $y\in\mathcal{Y}$, but it is easier to understand the meaning of the explicit results in the form “$k$-$l$”; i.e., the result $y_{t}=0$ corresponds to the outcome “3-0”, $y_{t}=1$ – to “3-1”, and $y_{t}=5$ – to “0-3”. We will use both and it should not lead to any confusion.
2.1 Ranking as statistical inference
The goal of the ranking is to order the teams and, to this end, we assume that they are characterized by intrinsic parameters called “skills” $\theta_{m}\in\mathbb{R},m=1,\ldots,M$. Then, by inferring the skills $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{M}]^{\mathsf{T}}$ from the observed outcomes of the matches $\boldsymbol{y}=[y_{1},\ldots,y_{T}]^{\mathsf{T}}$, and ordering them, the ranking is naturally obtained.
Arguably the most popular probabilistic model of the relationship between the skills $\boldsymbol{\theta}$, and the random variable $Y_{t}$ (which models $y_{t}$) links the latter to the difference between the skills of the home and away teams:
$\Pr\left\{Y_{t}=y|\boldsymbol{\theta},\boldsymbol{x}_{t}\right\}$ $=\mathsf{P}_{y}(z_{t}),$ (1)
$z_{t}$ $=\boldsymbol{x}_{t}^{\mathsf{T}}\boldsymbol{\theta},$ (2)
where $\mathsf{P}_{y}(z)$ is defined to strike balance between the complexity of the inference procedure and the modeling flexibility (that is, ability of the model (1) to fit the observations). Also, for compactness of notation, we introduce the scheduling vector
$\boldsymbol{x}_{t}=[x_{1,t},\ldots,x_{M,t}]^{\mathsf{T}},$ (3)
where $x_{i_{t},t}=1$, $x_{j_{t},t}=-1$, and $x_{m,t}=0$ for $m\notin\{i_{t},j_{t}\}$.
With the model defined, we may use conventional estimation strategies. For example, the maximum a posteriori (MAP) estimate of $\boldsymbol{\theta}$ is obtained solving
$\hat{\boldsymbol{\theta}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}\sum_{t=1}^{T}\ell_{y_{t}}(z_{t})+\rho(\boldsymbol{\theta}),$ (4)
where, the negated log-score
$\ell_{y}(z)$ $=-\log\mathsf{P}_{y}(z),$ (5)
and the prior distribution of the skills is defined via $\rho(\boldsymbol{\theta})=-\log f(\boldsymbol{\theta})$. The common assumption is to use the zero-mean Gaussian model for $\boldsymbol{\theta}$, and then
$\rho(\boldsymbol{\theta})$ $=\frac{1}{2}\gamma\|\boldsymbol{\theta}\|^{2}+{\textnormal{Const.}}$. (6)
A more general formulation is obtained by rewriting (4) as
$\hat{\boldsymbol{\theta}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}J(\boldsymbol{\theta})$ (7)
$J(\boldsymbol{\theta})$ $=\sum_{t=1}^{T}\ell^{{\textnormal{loss}}}_{y_{t}}(\boldsymbol{x}^{\mathsf{T}}_{t}\boldsymbol{\theta})+\rho(\boldsymbol{\theta}),$ (8)
where we use the “loss function” $\ell^{{\textnormal{loss}}}_{y}(z)$ which may, but does not need to, be the same as the log-score $\ell_{y}(z)$ in (5). As we shall see, to simplify calculations, the loss function may be a proxy of the log-score. In this “ordinal regression” formulation, $\rho(\boldsymbol{\theta})$ is called a regularization function, and using (6) yields the well-known L2 regularization.
Note that, using $\gamma=0$ and $\ell^{{\textnormal{loss}}}_{y}(z)=\ell_{y}(z)$ we obtain the conventional maximum likelihood (ML) estimation of the skills, which may be appropriate for “sufficiently large” $T$. However, with small/moderate $T$, it is prudent to use regularized form (7). [^3] Regularization is strictly necessary if there are teams whose matches finish with the extreme results “3-0” or “0-3” because then, without regularization, their skills tend to infinity.
Stochastic gradient ranking
We provide (7) to lay out a conceptual reference framework in which we can analyze the models used for ranking. This regression approach can be applied with a moderate value of $T$ (when the skills do not change significantly).
However, as also noted by Csató (2024), practical considerations of simplicity and transparency are more important in the context of sport ranking than the exact formulation. Thus, the online (real-time) ranking, where the skills are updated after each match, is more common in practice and is obtained by solving (7) using the stochastic gradient (SG) algorithm, defined as follows:
$\hat{\boldsymbol{\theta}}_{t+1}$ $=\hat{\boldsymbol{\theta}}_{t}-\mu\big{[}\nabla_{\boldsymbol{\theta}}\ell^{{\textnormal{loss}}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\boldsymbol{\theta})\big{]}_{\boldsymbol{\theta}=\hat{\boldsymbol{\theta}}_{t}}$ (9)
$=\hat{\boldsymbol{\theta}}_{t}-\mu\boldsymbol{x}_{t}\dot{\ell}^{{\textnormal{loss}}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}}_{t}),$ (10)
where $\mu$ is the adaptation step, and
˙$\dot{\ell}^{{\textnormal{loss}}}_{y}(z)$ $=\frac{\,\mathrm{d}}{\,\mathrm{d}z}\ell^{{\textnormal{loss}}}_{y}(z).$ (11)
The algorithm is initialized with $\boldsymbol{\theta}_{0}$, e.g., with $\boldsymbol{\theta}_{0}=[0,\ldots,0]^{\mathsf{T}}$.
When $\ell^{{\textnormal{loss}}}_{y}(z)$ is convex in $z$, for sufficiently large $t$ and appropriately chosen $\mu$, the solution of (10) approximates “well” $\hat{\boldsymbol{\theta}}$. The adaptation step, $\mu$, trades off the convergence speed (which tells how quickly, with $t$, $\hat{\boldsymbol{\theta}}_{t}$ approaches $\hat{\boldsymbol{\theta}}$) against the accuracy (which measures how far $\hat{\boldsymbol{\theta}}_{t}$ is from $\hat{\boldsymbol{\theta}}$). Although this is, admittedly, a vague statement, the precise analysis of the SG solution is not trivial even if some light is shed on this issue, e.g., in (Aldous, 2017), (Jabin and Junca, 2015), (Szczecinski and Roatis, 2022), (Gomes de Pinho Zanco et al., 2024). In fact, (10) may be seen as an approximation of a nonlinear Kalman filter, which estimates skills at time $t$ from previous observations $y_{1},\ldots,y_{t-1}$ (Szczecinski and Roatis, 2022, Sec. 3.3).
In this perspective, the skills $\hat{\boldsymbol{\theta}}_{t}$ are approximate solutions to the ML estimation of the skills $\boldsymbol{\theta}$: the approximation is due to the use of the stochastic gradient and due to the use of the loss function, which may approximate the log-score.
Scale
It turns out that the skills estimates, $\hat{\theta}_{m}$ ($\hat{\theta}_{m,t}$ in case of the SG algorithm (10)), may be quite small and therefore, to place them in a comfortable range (e.g., for visual interpretation by the users), we may multiply them by an arbitrarily chosen scale $s>0$, i.e., making the change of variables $\boldsymbol{\theta}^{\prime}=s\boldsymbol{\theta}$. This scale change can be made directly on the final solutions in (7), as $\hat{\boldsymbol{\theta}}^{\prime}=s\hat{\boldsymbol{\theta}}$ or in (10), as $\hat{\boldsymbol{\theta}}^{\prime}_{t}=s\hat{\boldsymbol{\theta}}_{t}$. In fact, the latter can be integrated into the recursive equation yielding
$\hat{\boldsymbol{\theta}}^{\prime}_{t+1}$ $=\hat{\boldsymbol{\theta}}^{\prime}_{t}-\mu s\boldsymbol{x}_{t}\dot{\ell}^{{\textnormal{loss}}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}}^{\prime}_{t}/s).$ (12)
2.2 Integrating exogenous variables
We assumed that the probabilistic model (1) depends only on $z_{t}=\boldsymbol{x}^{\mathsf{T}}_{t}\boldsymbol{\theta}$ and the outcome $y_{t}$. However, a more general approach may be used in which other exogenous variables affect the model.
Home-field advantage (HFA)
For example, we may want to take into account the HFA using a binary variable $h_{t}\in\{0,1\}$ that indicates whether the match $t$ is played on the home venue, i.e., in the country of the team $i_{t}$ (then $h_{t}=1$), or is played in the neutral venue (then $h_{t}=0$). The popular model relies on boosting the skills of the home team, or, equivalently, on increasing the values of $z_{t}$ by the HFA parameter $\eta$, i.e.,
$\ell^{{\textnormal{loss}}}_{y_{t},h_{t}}(z_{t})=\ell^{{\textnormal{loss}}}_{y_{t}}(z_{t}+h_{t}\eta);$ (13)
the exogenous variable $h_{t}$ is shown as a subscript, and $\eta$ is a part of the model.
Weighting/matches importance
Using the same notation, we may modify the loss function via heuristics, such as weighting
$\ell^{{\textnormal{loss}}}_{y_{t},v_{t}}(z_{t})$ $=\xi_{v_{t}}\ell^{{\textnormal{loss}}}_{y_{t}}(z_{t})$ (14)
where $v_{t}$ is a categorical variable associated with match $t$, $v_{t}\in\{0,\ldots,K-1\}$, and the weight $\xi_{v_{t}}\geq 0$ allows us to modulate the relative importance of the term $\ell^{{\textnormal{loss}}}_{y}(z)$: the smaller $\xi_{v_{t}}$ is, the less impact the pair $(z_{t},y_{t})$ will have on the solution $\hat{\boldsymbol{\theta}}$.
In the ranking context, $v_{t}$ is called the “prestige” of the match (in the FIVB ranking) or its “importance” (in the FIFA ranking).
[…]
2.3 Current FIVB ranking algorithm
The FIVB ranking defines (1)
$\mathsf{P}^{{\textnormal{FIVB}}}_{y}(z)$ $=\mathsf{P}_{y}(z;\boldsymbol{c}^{{\textnormal{FIVB}}})$ (15)
$\mathsf{P}_{y}(z;\boldsymbol{c})$ $=\Phi(z+c_{y})-\Phi(z+c_{y-1}),\quad y=0,\ldots,L-1,$ (16)
where $\Phi(z)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{z}\exp(-0.5v^{2})\,\mathrm{d}v$ is the cumulative density function (CDF) of a zero-mean, unit-variance Gaussian distribution.
The model (16) is an extension of the Bradley-Terry model for binary outcomes (Bradley and Terry, 1952) and was also used with ternary outcomes by Glenn and David (1960). The generalization to multi-level outcomes in (16) is known as the ordinal probit (Agresti, 2013, Ch. 8.3); we call it a “CL model” (Tutz, 2012, Ch. 9.1) which is more general allowing us to use non-Gaussian CDF s in the model.
The model is parameterized with thresholds $\boldsymbol{c}=[c_{-1},c_{0},\ldots,c_{L-1}]$, which are monotonically increasing with $l$, and, to simplify the discussion, may also be assumed symmetric, i.e.,
$c_{l}=-c_{L-2-l},$ (17)
and we always set the first, and the last thresholds as $c_{-1}=-c_{L-1}=-\infty$. Also, for even $L$, we have $c_{\frac{L}{2}-1}\equiv 0$. In particular, in the FIVB model, with $L=6$, we always have $c_{2}\equiv 0$, and $-c_{-1}=c_{5}=\infty$ and thus only two parameters can be set independently: $c_{0}$ and $c_{1}$.
The FIVB ranking defines the thresholds as follows:
$c^{{\textnormal{FIVB}}}_{0}=-c^{{\textnormal{FIVB}}}_{4}=-1.06,$ $c^{{\textnormal{FIVB}}}_{1}=-c^{{\textnormal{FIVB}}}_{3}=-0.394,$ $c^{{\textnormal{FIVB}}}_{2}=0.$ (18)
The symmetry of the thresholds $c^{{\textnormal{FIVB}}}_{y}=-c^{{\textnormal{FIVB}}}_{L-2-y}$, and of the CDF, $\Phi(z)=1-\Phi(-z)$, yields the symmetric forms $\mathsf{P}^{{\textnormal{FIVB}}}_{y}(z)=\mathsf{P}^{{\textnormal{FIVB}}}_{L-1-y}(-z)$, as can be appreciated in Fig. 1. One of the properties of the CL model is that the probability of the outcome “not less important than $y$” is calculated as
$\Pr\left\{Y\leq y|z\right\}=\sum_{l=0}^{y}\mathsf{P}^{{\textnormal{FIVB}}}_{l}(z)=\Phi(z+c^{{\textnormal{FIVB}}}_{y}).$ (19)
Figure 1: Functions $\mathsf{P}^{{\textnormal{FIVB}}}_{y}(z)$ in the FIVB ranking algorithm, where $y=0$ corresponds to the result “3-0”, $y=1$ to “3-1”, etc.
With the model defined, the FIVB ranking algorithm (FIVB, 2024) estimates the skills as follows:
$\boldsymbol{\theta}_{t+1}$ $=\boldsymbol{\theta}_{t}-\mu s\xi_{v_{t}}\boldsymbol{x}_{t}g^{{\textnormal{FIVB}}}_{y_{t}}(z_{t}/s),$ (20)
$g^{\textnormal{FIVB}}_{y}(z)$ $=\check{r}(z)-r^{\textnormal{FIVB}}_{y},$ (21)
where the adaptation step is $\mu=0.01$, the scale is $s=125$, the weights $\xi_{v}$ are defined in Table 1, and $r^{\textnormal{FIVB}}_{y}$ is the numerical score assigned to the outcome $y$, see Table 2. It may be surprising because, as we emphasized previously, ordinal outcomes have no numerical value. This is still true: in fact, as we will see in Sec. 2.5, variables $r^{\textnormal{FIVB}}_{y}$ are merely auxiliary parameters defining the loss function $\ell^{{\textnormal{loss}}}_{y}(z)$, which is not the same as the log-score (5).
The probabilistic model (16) is used to calculate the expected value of $r^{\textnormal{FIVB}}_{Y}$ (for a given $z$):
$\check{r}(z)$ $=\mathds{E}_{Y|z}[r^{\textnormal{FIVB}}_{Y}]=\sum_{y\in\mathcal{Y}}r^{\textnormal{FIVB}}_{y}\mathsf{P}^{{\textnormal{FIVB}}}_{y}(z)$ (22)
$=\sum_{y=0}^{L-2}(r^{\textnormal{FIVB}}_{y}-r^{\textnormal{FIVB}}_{y+1})\Phi(z+c^{{\textnormal{FIVB}}}_{y})+r^{\textnormal{FIVB}}_{L-1}.$ (23)
$v$ $\xi_{v}$ Description
00 $1.00$ Official events of Continental Confederations
$1$ $1.75$ Confederations’ Championship qualifying
$2$ $2.00$ FIVB Challenger Cup
$3$ $3.50$ Olympic matches qualifying,
Confederations’ Championship
$4$ $4.00$ FIVB Nations League
$5$ $4.50$ FIVB World Championship
$6$ $5.00$ Olympic matches
Table 1: The weighting coefficients adopted from the FIVB ranking. Note that we normalize $\xi_{v}$ by using $\xi_{v}=1.0$.
$y$ “3-0” “3-1” “3-2” “2-3” “1-3” “0-3”
$r^{\textnormal{FIVB}}_{y}$ 2.0 1.5 1.0 -1.0 -1.5 -2.0
Table 2: Numerical scores $r^{\textnormal{FIVB}}_{y}$ assigned to the outcomes $y$ in the FIVB ranking algorithm.
The FIVB algorithm defined by (20)-(22) is meant to be compatible with the formulation in (12) but, of course, the official FIVB presentation (FIVB, 2024) uses a different notation. For example, instead of defining the step, the scale and the weights, the FIVB ranking defines their product $\mu s\xi_{v}$, which is required in (20); we explain this in detail in Appendix A.
The FIVB ranking has an additional rule, where, at the end of the year, each team $m$ that did not participate in any competition has the value of its skills reduced $\theta_{m,t}\leftarrow\theta_{m,t}-50$. The objective is to discourage match avoidance to preserve the value of the skills (and thus the ranking position). But these are heuristics that cannot be easily put in the statistical framework, and thus we do not model them.
2.4 Data
To evaluate the FIVB ranking algorithm, we use the matches played by the men’s national teams and published on the FIVB website (FIVB, 2024). We only consider the matches since January 1, 2021, as, before that date, due to the Covid-19 pandemic, many matches were eliminated and some had the weighting $\xi_{v}$ incompatible with the official description. [^5] For example, the matches played in January 2020 used the weight $\xi_{v}=2.50$ which is not specified in the ranking. The results were collected till the end of December 2023 (with no matches after October 2023).
Since the FIVB website publishes the result $y_{t}$ of the match $t$ and the change in the value of skills, $\theta_{m,t+1}-\theta_{m,t}$, we can find the value of $\xi_{v_{t}}$ and infer the match category $v_{t}$ from Table 1. To establish the venue of the match (which affects $h_{t}$), we relied on Wikipedia pages describing the international volleyball events that we paired with matches from the FIVB website. [^6] These results were non exhaustively validated by comparing with the venues of main international events shown on the FIVB website (such as Nations League, Challenger Cup or World Championship).
Moreover, we remove
  • The matches in which, the FIVB ranking displays increments with small absolute value $|\theta_{m,t+1}-\theta_{m,t}|\in\{0,0.01\}$. [^7] There were $67$ matches with the absolute value of the increment equal to $0.01$, and $33$ with the increment equal to $0.0$. For example, the Oct. 8, 2023 match China-Poland had the increment equal to $0.01$, and the 24 Sept., 2023, USA-Canada match had the increment equal to $0.0$. This may happen, e.g., when players could not obtain visas, and sometimes it is explained on the FIVB website (e.g., in the case of the Iranian team playing in the USA in 2023) but, in other cases, the non-standard values of $\xi$ are left unexplained. Since these matches do not contribute to the change in the ranking, excluding them from considerations is consistent with the spirit of the FIVB ranking.
  • The forfeited matches we could identify, [^8] Including: i) Denmark’s matches in Jan. 2021, ii) Uzbekistan’s and Pakistan’s matches in July 2023, and iii) Mongolia’s matches in Aug. 2023. Even if FIVB treats a forfeited match as a “$3-0$” win, in our opinion, the match which did not take place does not reflect on the strength of the team, and thus we do not use it as information for ranking.
Then, we have a total of $M=102$ teams and $T=1151$ matches. To characterize the results, we count matches with result $y_{t}=y$, played on the neutral- and home-venues
$k^{{\textnormal{ntr}}}_{y}$ $=\frac{1}{2}\sum_{t=1}^{T}\big{(}\mathds{I}\big{[}{y_{t}=y,h_{t}=0}\big{]}+\mathds{I}\big{[}{y_{t}=L-1-y,h_{t}=0}\big{]}\big{)},$ (24)
$k^{{\textnormal{hfa}}}_{y}$ $=\sum_{t=1}^{T}\mathds{I}\big{[}{y_{t}=y,h_{t}=1}\big{]}$ (25)
and show them in Table 3.
Note that in neutral-venue matches, there is no distinction between the home/away teams, so the number of results “$k-l$” (denoted as $y$) and “$l-k$” (denoted as $L-1-y$) should be equal: this is why we count all these results in (24). Although this produces a fractional value $k^{{\textnormal{ntr}}}_{y}$, this formalism simplifies the notation, and we guarantee that $k^{{\textnormal{ntr}}}_{y}=k^{{\textnormal{ntr}}}_{L-1-y}$.
$y$ “3-0” “3-2” “3-1” “1-3” “2-3” “0-3” Total
$k^{{\textnormal{ntr}}}_{y}$ 203.5 117.5 59.5 59.5 117.5 203.5 761
$k^{{\textnormal{hfa}}}_{y}$ 135 64 29 33 45 84 390
Table 3: Numbers of the FIVB matches with outcomes $y\in\mathcal{Y}$, played on the neutral- and home- venues, and their totals.
2.5 Implicit loss function
Knowing the probabilistic model, we can derive the SG algorithm setting $\ell_{y}^{{\textnormal{loss}}}(z)=\ell_{y}(z)$, i.e., using the log-score as a loss function with
$\ell_{y}(z)$ $=-\log\big{(}\Phi(z+c_{y})-\Phi(z+c_{y-1})\big{)},$ (26)
and finding the derivative of the latter
˙$\dot{\ell}_{y}(z)$ $=-\frac{\mathcal{N}(z+c_{y})-\mathcal{N}(z+c_{y-1})}{\Phi(z+c_{y})-\Phi(z+c_{y-1})},$ (27)
where we use the Gaussian probability density function (PDF)
$\mathcal{N}(z)=\dot{\Phi}(z)=\frac{1}{\sqrt{2\pi}}\exp(-0.5z^{2}).$ (28)
Quite obviously, plugging ˙$\dot{\ell}_{y}(z)$ shown in (27), into the SG algorithm (12), does not produce the FIVB ranking (20). Thus, the FIVB ranking does not solve the ML or MAP problem.
To understand what problem it does solve, we note that, since the FIVB ranking (20) has the structure of the SG optimization (12), we may treat the function $g^{{\textnormal{FIVB}}}_{y}(z)$ used by (20) as a derivative of an implicit (that is, not explicitly defined) loss function $\ell^{{\textnormal{FIVB}}}_{y}(z)$, i.e., $g^{{\textnormal{FIVB}}}_{y}(z)=\dot{\ell}^{{\textnormal{FIVB}}}_{y}(z)$, and, the latter can be unveiled through the integration of $g^{{\textnormal{FIVB}}}_{y}(z)$, i.e.,
$\ell^{{\textnormal{FIVB}}}_{y}(z)$ $=\int_{-\infty}^{z}g^{{\textnormal{FIVB}}}_{y}(u)\,\mathrm{d}u=\int_{-\infty}^{z}\big{[}\check{r}(u)-r^{\textnormal{FIVB}}_{y}\big{]}\,\mathrm{d}u,$ (29)
$=\sum_{l=0}^{L-2}(r^{{\textnormal{FIVB}}}_{l}-r^{{\textnormal{FIVB}}}_{l+1})\psi(z+c^{{\textnormal{FIVB}}}_{l})+(r^{\textnormal{FIVB}}_{L-1}-r^{\textnormal{FIVB}}_{y})z+{\textnormal{Const.}}$. (30)
where (30) is obtained from (23) and
$\psi(z)$ $=\int_{-\infty}^{z}\Phi(u)\,\mathrm{d}u=\Phi(z)z+\mathcal{N}(z).$ (31)
To understand why the FIVB ranking algorithm does not use directly the log-score (26), two issues with (27) should be considered.
The first is of numerical nature, because, for large $|z|$, both the numerator and the denominator in (27) tend to zero which requires a careful implementation.
The second problem is that (27) has a relatively complicated form without obvious interpretation. In that regard, (21) has a practical advantage: the skills $\boldsymbol{\theta}_{t}$ are updated using the difference between the observed and expected numerical scores. Note that this is also the usual interpretation of the well-known Elo algorithm (Elo, 2008).
We therefore conjecture that the FIVB ranking algorithm was designed to be simple and understandable, which is a typical requirement in the sport ranking (Csató, 2024, Sec. 1). The potential drawback is the sub-optimality of the ranking results, which we will evaluate in this work.
Convergence
A minor point is to obtain a guarantee that, for a sufficiently small $\mu$, the FIVB algorithm (20) converges, for which we need the following:
Lemma 1.
The implicit loss function, $\ell^{{\textnormal{FIVB}}}_{y}(z)$ is convex in $z$.
Proof:
The second derivative of implicit loss $\ell^{{\textnormal{FIVB}}}_{y}(z)$ is positive if the first derivative $g^{{\textnormal{FIVB}}}_{y}(z)=\check{r}(z)-r^{\textnormal{FIVB}}_{y}$ increases monotonically in $z$. The latter holds because (23) is monotonically increasing if $r^{\textnormal{FIVB}}_{y+1}<r^{\textnormal{FIVB}}_{y}$, which is true from Table 2.
We note that condition $r^{\textnormal{FIVB}}_{y+1}<r^{\textnormal{FIVB}}_{y}$ is sufficient but not necessary, i.e., it can be violated, and yet we can still obtain a convex function $\ell^{{\textnormal{FIVB}}}_{y}(z)$ (see examples in Sec. 3.3.2). The fact that this is possible only reinforces the idea that numerical scores are auxiliary parameters and should not be thought of as a value of the match outcome (because the latter simply do not exist).
3 Model identification
In this part of the work, we assume that
  • The probabilistic model (16) is predefined, and we need to define the thresholds $\boldsymbol{c}$ and, eventually, the HFA coefficient $\eta$.
  • The loss function is defined via $\ell^{{\textnormal{FIVB}}}_{y}(z)$ in (30), and we need to define the suitable numerical scores $\boldsymbol{r}=[r_{0},\ldots,r_{L-1}]$ which are attributed to the outcomes, and
  • The weights $\boldsymbol{\xi}$ are used in the algorithm via (14), and we want to assess their usefulness.
In order to assess and/or optimize the parameters of the model, we need a well-defined criterion.
3.1 Model identification via cross-validation
Inference (7) may be done if we define (the form of) the loss and regularization functions, as well as, if we find their parameters $\boldsymbol{p}$ (also called hyper-parameters, in the machine learning language) that affect all the functions describing the models, i.e., we can write $\ell^{{\textnormal{loss}}}(z)\equiv\ell^{{\textnormal{loss}}}_{y}(z;\boldsymbol{p})$, $\ell_{y}(z)\equiv\ell_{y}(z;\boldsymbol{p})$ and $\rho(\boldsymbol{\theta})\equiv\rho(\boldsymbol{\theta};\boldsymbol{p})$.
A well-known approach to finding the hyper-parameters, particularly suitable for relatively small data sets, relies on leave-one-out (LOO) cross-validation (Hastie et al., 2009, Ch. 2.9), where we remove one observation $y_{t}$, and we verify how well the results (7), denoted as $\hat{\boldsymbol{\theta}}_{\backslash{t}}$, match the removed data using a metric $\ell^{{\textnormal{val}}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}}_{\backslash{t}})$; for the latter, most often we use the log-score $\ell^{{\textnormal{val}}}_{y}(z)=\ell_{y}(z)$. This is repeated for all $T$ samples and may be defined as follows (Hastie et al., 2009, Ch. 2.9)
$\hat{\boldsymbol{p}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{p}}U(\boldsymbol{p})$ (32)
$U(\boldsymbol{p})$ $=\frac{1}{T}\sum_{t=1}^{T}\ell_{y_{t}}^{{\textnormal{val}}}(\hat{z}_{t,\backslash{t}};\boldsymbol{p})$ (33)
$\hat{z}_{t,\backslash{t}}$ $=\boldsymbol{x}^{\mathsf{T}}_{t}\hat{\boldsymbol{\theta}}_{\backslash{t}}$ (34)
$\hat{\boldsymbol{\theta}}_{\backslash{t}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}\big{[}J(\boldsymbol{\theta})-\ell^{{\textnormal{loss}}}_{y_{t}}(\boldsymbol{x}^{\mathsf{T}}_{t}\boldsymbol{\theta};\boldsymbol{p})\big{]}.$ (35)
where $U(\boldsymbol{p})$ is the averaged validation metric. [^9] Note that, while the LOO cross-validation uses the validation sets containing only one element, in general, we can use larger sets. However, they must then be defined arbitrarily (e.g., randomly), as it is rarely possible to enumerate all of them. On the other hand, in the LOO approach we do enumerate all $T$ validation sets. Thus, given the data, the results are independent of the random/arbitrary definition of the validation sets. This removes any ambiguity in the numerical optimization of $U(\boldsymbol{p})$ required in this work.
We can also calculate the metrics for the matches played on the neutral and the home venues, given, respectively, by
$U^{{\textnormal{ntr}}}(\boldsymbol{p})$ $=\frac{1}{T^{{\textnormal{ntr}}}}\sum_{\begin{subarray}{c}t=1\\ h_{t}=0\end{subarray}}^{T}\ell^{{\textnormal{val}}}(\hat{z}_{t,\backslash{t}};\boldsymbol{p})$ (36)
$U^{{\textnormal{hfa}}}(\boldsymbol{p})$ $=\frac{1}{T^{{\textnormal{hfa}}}}\sum_{\begin{subarray}{c}t=1\\ h_{t}=1\end{subarray}}^{T}\ell^{{\textnormal{val}}}(\hat{z}_{t,\backslash{t}};\boldsymbol{p}),$ (37)
where, $U(\boldsymbol{p})=\big{(}T^{{\textnormal{ntr}}}U^{{\textnormal{ntr}}}(\boldsymbol{p})+T^{{\textnormal{hfa}}}U^{{\textnormal{hfa}}}(\boldsymbol{p})\big{)}/T$.
We note that (33) can be transformed as follows:
$V(\boldsymbol{p})=\mathrm{e}^{-U(\boldsymbol{p})}=\Big{[}\prod_{t=1}^{T}\mathsf{P}_{y_{t}}(\hat{z}_{t,\backslash{t}};\boldsymbol{p}))\Big{]}^{\frac{1}{T}},$ (38)
which is the geometric mean of the probabilities of observing $Y_{t}=y_{t}$, calculated from $\hat{\boldsymbol{\theta}}_{\backslash{t}}$. Thus, $V(\boldsymbol{p})$ is potentially easier to interpret than $U(\boldsymbol{p})$; and, with linearization, for small $U(\boldsymbol{p})$, e.g., $U(\boldsymbol{p})<0.1$, we have $V(\boldsymbol{p})\approx 1-U(\boldsymbol{p})$. For example, $U(\boldsymbol{p})=1.4$ yields $V(\boldsymbol{p})\approx 24.7$%, while for a uniform distribution $\mathsf{P}_{y}(z)=\frac{1}{6},\forall y$ (i.e., $V(\boldsymbol{p})=16.7$%) we have $U(\boldsymbol{p})=1.79$. The values $V(\boldsymbol{p})$ are shown on the right-hand auxiliary axis in Figs. 3, Fig. 5, and Fig. 7.
While the complexity is reduced, we still need to solve (35) $T$ times, and, to alleviate the complexity, we apply here the approximate leave-one-out (ALO) approach (Beirami et al., 2017),(Rad and Maleki, 2020), (Burn, 2020), where (7) is solved once to find $\hat{\boldsymbol{\theta}}\equiv\hat{\boldsymbol{\theta}}(\boldsymbol{p})$ (this is where most of the computational complexity lies), and we make a quadratic approximation of the function $J(\boldsymbol{\theta})$ around $\hat{\boldsymbol{\theta}}$, which allows us to find the closed-form approximation of (34) as follows:
$\hat{z}_{t,\backslash{t}}$ $\approx\hat{z}_{t}+\frac{\dot{\ell}^{{\textnormal{loss}}}_{y_{t}}(\hat{z}_{t};\boldsymbol{p})a_{t}}{1-\ddot{\ell}^{{\textnormal{loss}}}_{y_{t}}(\hat{z}_{t};\boldsymbol{p})a_{t}},$ (39)
where $a_{t}=\boldsymbol{x}^{\mathsf{T}}_{t}\hat{{\textnormal{{H}}}}^{-1}\boldsymbol{x}_{t}$, $\hat{{\textnormal{{H}}}}=\nabla^{2}_{\boldsymbol{\theta}}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\hat{\boldsymbol{\theta}}}$ is the Hessian of the function $J(\boldsymbol{\theta})$ under minimization in (7), ¨$\ddot{\ell}^{{\textnormal{loss}}}_{y}(z;\boldsymbol{p})=\frac{\,\mathrm{d}}{\,\mathrm{d}z}\dot{\ell}^{{\textnormal{loss}}}_{y}(z;\boldsymbol{p})$, and $\hat{z}_{t}=\boldsymbol{x}^{\mathsf{T}}_{t}\hat{\boldsymbol{\theta}}(\boldsymbol{p})$. Details of the derivation can be found in (Burn, 2020, Sec. 3) or (Szczecinski and Roatis, 2022, Appendix. 2).
3.2 Finding thresholds $c$ and HFA parameter $\eta$
To assess the role of the thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$ used in the CL model (16) and of the HFA parameter $\eta$, we start with $\ell^{{\textnormal{loss}}}_{y}(z)=\ell_{y}(z)$ and consider four cases:
  1. i)
    We use $\boldsymbol{c}=\boldsymbol{c}^{{\textnormal{FIVB}}}$ given by (18) and $\eta=0$; in other words, no optimization is performed.
  2. ii)
    We find $\hat{\boldsymbol{c}}$ by optimizing $U(\boldsymbol{p})$ in (32) but we set $\eta=0$, i.e.,
[…]
  3. iii)
    We use $\boldsymbol{c}^{{\textnormal{FIVB}}}$ and optimize the HFA parameter, i.e., use
    $\hat{\eta}=\mathop{\mathrm{argmin}}_{\eta,\boldsymbol{c}=\boldsymbol{c}^{{\textnormal{FIVB}}}}U(\eta);$ (41)
  4. iv)
    We find both $\hat{\boldsymbol{c}}$ and $\hat{\eta}$ through optimization,
[…]
    and they are shown in Fig. 2.
In the above, we slightly abuse the notation and write e.g., $U(\boldsymbol{c})$ to indicate that we optimize only the parameter $\boldsymbol{c}$ and all other hyperparameters in $\boldsymbol{p}$ are kept constant (as we also explicitly indicate, when relevant, under the $\mathop{\mathrm{argmin}}$ operator); similarly, the notation $U(\boldsymbol{c},\eta)$ means that both $\boldsymbol{c}$ and $\eta$ are optimized.
Figure 2: Results of optimization in (42): thresholds $\hat{c}_{0}$, $\hat{c}_{1}$ and the HFA $\hat{\eta}$. The horizontal dotted lines are drawn for the values $c^{{\textnormal{FIVB}}}_{0}=-1.06$ and $c^{{\textnormal{FIVB}}}_{1}=-0.394$ used by the FIVB ranking, see (18). All thresholds satisfy (17) so $c_{2}\equiv 0$ need not be shown.
We show, in Fig. 3, cross-validation results $U^{{\textnormal{ntr}}}(\boldsymbol{p})$ and $U^{{\textnormal{hfa}}}(\boldsymbol{p})$ defined in (36) and (37) as functions of the regularization parameter $\gamma$, where $\boldsymbol{p}$ contains all parameters, including $\gamma$ and others that are optimized according to the cases we explain above. The right-hand axis in Fig. 3 shows the metric $V(\boldsymbol{p})$ as a transformation of $U(\boldsymbol{p})$ through (38).
Figure 3: Validation metrics $U^{{\textnormal{ntr}}}$ (hollow markers) and $U^{{\textnormal{hfa}}}$ (colored markers) given by (36)-(37) shown as a function of $\gamma$ using the loss function $\ell_{y}(z)$ defined in (5).
Main observations
From Fig. 2, we see that the FIVB thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$ are close to, but not identical with those obtained through optimization. This discrepancy should be attributed to the difference in the data sets from which the thresholds were inferred, but also to the procedure described by the FIVB ranking, which defined $\boldsymbol{c}^{{\textnormal{FIVB}}}$ from the matches of the “teams with similar skills”. [^10] We note that this approach to find the model is rather ambiguous as, to find which skills are similar, we have to estimate them, and this requires the model to be defined in the first place. In our approach, we used all matches, which may explain some improvement in the validation metrics $U(\hat{\boldsymbol{p}})$ we observe in Fig. 3.
From Fig. 3 we observe that:
  • The decomposition into the matches played on the neutral and home venues indicates that, by including the HFA into the model via $\eta$, we can improve the prediction for the home matches, and, rather unsurprisingly, this modification has practically no impact on the neutral-venue matches.
  • Attention should be paid to the metric $V(\boldsymbol{p})$ shown on the right axis, where we see that the improvements are on the order of a fraction of a percent. For example, the most significant improvement appears for home matches, where, by optimizing $\eta$ and the thresholds $\boldsymbol{c}$ in (42), the results are improved by $\sim 1\%$; i.e., from $V\approx 25.8\%$ to $V\approx 26.8\%$.
3.3 Finding numerical score $r$
To discuss the role of the numerical scores $\boldsymbol{r}$ used in the FIVB algorithm, we will first, in Sec. 3.3.1, analyze the implicit loss functions used by the algorithm while a purely numerical analysis / optimization is carried out in Sec. 3.3.2.
3.3.1 From thresholds to numerical scores
We show, in Fig. 4, the logarithmic loss functions $\ell_{y}(z)$, as well as, the scaled and vertically-shifted versions of the implicit loss functions $a\ell^{{\textnormal{FIVB}}}_{y}(z;\boldsymbol{r})+b_{y}$, where we find $a$ by matching the first derivatives of the loss function $\ell_{0}(z_{{\textnormal{o}}})$ for $z=z_{{\textnormal{o}}}=0$
˙$\dot{\ell}_{0}(z_{{\textnormal{o}}})$ $=ag^{{\textnormal{FIVB}}}_{0}(z_{{\textnormal{o}}};\boldsymbol{r}),$ (43)
which, from $g^{{\textnormal{FIVB}}}_{y}(z_{{\textnormal{o}}};\boldsymbol{r})=\check{r}(z_{{\textnormal{o}}})-r_{y}=-r_{y}$, [^11] Because, due to symmetry, the expected numerical score is zero, i.e., $\check{r}(0)=0$ yields
$a$ =˙$=-\frac{\dot{\ell}_{0}(z_{{\textnormal{o}}})}{r_{0}}.$ (44)
Similarly, the shifts $b_{y}$ are calculated to match the values of the loss functions at $z_{{\textnormal{o}}}$, i.e., to satisfy $\ell_{y}(z_{{\textnormal{o}}})=a\ell^{{\textnormal{FIVB}}}_{y}(z_{{\textnormal{o}}};\boldsymbol{r})+b_{y}$. This means that $b_{y}=\ell_{y}(z_{{\textnormal{o}}})-a\ell^{{\textnormal{FIVB}}}_{y}(z_{{\textnormal{o}}};\boldsymbol{r})$.
This scaling/shifting transformation is irrelevant from the optimization point of view [^12] Scaling all loss function with $a>0$, obviously, does not change the results of (4). Similarly, adding $b_{y}$ to each loss function does not affect optimality, but allows us to visually appreciate the difference between the loss functions in the vicinity of the target value $z_{{\textnormal{o}}}=0$. Note that assuming that $z_{t}$ will be mostly observed close to $z_{{\textnormal{o}}}$ is compatible with $z$ being a zero-mean Gaussian variable, which is the modeling assumption we used in Sec. 2.1.
Figure 4: Loss functions: log-score $\ell_{y}(z)$ (solid line) and rescaled/shifted implicit loss function $a\ell^{{\textnormal{FIVB}}}_{y}(z;\boldsymbol{r})+b_{y}$ (dashed line) for $y=0,1,2$ corresponding, respectively, to the outcomes “3-0”, “3-1”, and “3-2”.
Indeed, Fig. 4 indicates that an almost perfect match is obtained for $\ell_{0}(z)$, i.e., the logarithmic loss is practically indistinguishable from the scaled/shifted implicit loss. However, the loss functions are not matched for $y=1,2$, and we want to find the numerical score $\tilde{\boldsymbol{r}}=[\tilde{r}_{0},\ldots,\tilde{r}_{L-1}]$, which satisfies a generalized version of (43), i.e., we want the latter to hold for all $y\in\mathcal{Y}$, that is
˙$\dot{\ell}_{y}(z_{{\textnormal{o}}})=ag_{y}^{{\textnormal{FIVB}}}(z_{{\textnormal{o}}};\tilde{\boldsymbol{r}}),\quad y\in\mathcal{Y}.$ (45)
Since ˙$\dot{\ell}_{y}(z_{0})=g_{y}^{{\textnormal{FIVB}}}(z_{{\textnormal{o}}};\tilde{\boldsymbol{r}})=-\tilde{r}_{y}$, (45) is solved by
$\tilde{r}_{y}\equiv\tilde{r}_{y}(\boldsymbol{c})$ $=-\tilde{r}_{0}\frac{\dot{\ell}_{y}(z_{{\textnormal{o}}})}{\dot{\ell}_{0}(z_{{\textnormal{o}}})}=\tilde{r}_{0}\frac{\Phi(c_{0})\big{(}\mathcal{N}(c_{y})-\mathcal{N}(c_{y-1})\big{)}}{\mathcal{N}(c_{0})\big{(}\Phi(c_{y})-\Phi(c_{y-1})\big{)}},\quad y\in\mathcal{Y},$ (46)
where we used (27) and (44); the notation $\tilde{r}_{y}(\boldsymbol{c})$ emphasizes the dependence of $\tilde{\boldsymbol{r}}$ on the thresholds $\boldsymbol{c}$ which define the CL model so (46) is valid for any $\boldsymbol{c}$. Note also that $\tilde{r}_{y}$ is not uniquely defined because we can arbitrarily fix $\tilde{r}_{0}$, and, for comparison with the FIVB ranking, we set $\tilde{r}_{0}\equiv r^{{\textnormal{FIVB}}}_{0}=2.0$.
We can now calculate the scores by applying (46) to the FIVB-defined thresholds, $\boldsymbol{c}=\boldsymbol{c}^{{\textnormal{FIVB}}}$, which yields $\tilde{\boldsymbol{r}}=[\tilde{r}_{0}(\boldsymbol{c}^{{\textnormal{FIVB}}}),\ldots,\tilde{r}_{L-1}(\boldsymbol{c}^{{\textnormal{FIVB}}})]$
$\tilde{r}_{0}$ $=-\tilde{r}_{5}\equiv 2.0,$ (47)
$\tilde{r}_{1}(\boldsymbol{c}^{{\textnormal{FIVB}}})$ $=-\tilde{r}_{4}(\boldsymbol{c}^{{\textnormal{FIVB}}})=0.89,$ (48)
$\tilde{r}_{2}(\boldsymbol{c}^{{\textnormal{FIVB}}})$ $=-\tilde{r}_{3}(\boldsymbol{c}^{{\textnormal{FIVB}}})=0.25.$ (49)
These values are rather different from those shown in Table 2 and, more importantly, when we use them to calculate the implicit loss functions $\ell^{{\textnormal{FIVB}}}_{y}(z;\tilde{\boldsymbol{r}}),y=0,1,2$ (shown, scaled with $\tilde{a}$ and shifted with $\tilde{b}_{y}$, in Fig. 4 with dashed-dotted lines), the latter are indistinguishable from $\ell_{y}(z)$ in the vicinity of $z_{{\textnormal{o}}}=0$. Clearly, when compared to the implicit loss $\ell_{y}^{{\textnormal{FIVB}}}(z;\boldsymbol{r}^{{\textnormal{FIVB}}})$ with FIVB-defined scores $\boldsymbol{r}^{{\textnormal{FIVB}}}$, the loss $\ell^{\textnormal{FIVB}}_{y}(z;\tilde{\boldsymbol{r}})$ offers a better approximation of the log-loss $\ell_{y}(z)$. And this improvement is obtained solely using the numerical score $\tilde{\boldsymbol{r}}$.
3.3.2 Numerical optimization
Instead of the analytical approach, shown in the previous section, the numerical optimization of $\boldsymbol{r}$, takes into account the actual outcomes of the matches.
To analyze the optimality of $\boldsymbol{r}^{{\textnormal{FIVB}}}$, the loss function $\ell^{{\textnormal{loss}}}_{y}(z)$ is set to $\ell^{{\textnormal{FIVB}}}_{y}(z)$, and we consider the following cases:
  1. i)
    We use $\boldsymbol{c}^{\textnormal{FIVB}}$ and $\boldsymbol{r}^{\textnormal{FIVB}}$ specified by the FIVB ranking, and given, respectively, in (18) and Table 2, and we set $\eta=0$. This is the reference, currently used by FIVB.
  2. ii)
    We optimize $\boldsymbol{r}$, $\eta$ for a given $\boldsymbol{c}$
    $\hat{\boldsymbol{r}}(\boldsymbol{c}),\hat{\eta}(\boldsymbol{c})=\mathop{\mathrm{argmin}}_{\boldsymbol{r},\eta}U(\boldsymbol{r},\eta,\boldsymbol{c},\boldsymbol{p}_{\backslash\{\boldsymbol{r},\eta,\boldsymbol{c}\}}),$ (50)
    where we consider $\boldsymbol{c}=\boldsymbol{c}^{\textnormal{FIVB}}$ and $\boldsymbol{c}=\hat{\boldsymbol{c}}$, with the latter obtained via (42).
  3. iii)
    We calculate the numerical score $\tilde{\boldsymbol{r}}=\tilde{\boldsymbol{r}}\big{(}\boldsymbol{c}\big{)}$ using the formula shown in (46), and set $\eta=0.2$ (which is a rounded value of $\hat{\eta}$ obtained through optimization; see Fig. 2). As before, this is done for $\boldsymbol{c}=\boldsymbol{c}^{\textnormal{FIVB}}$ and $\boldsymbol{c}=\hat{\boldsymbol{c}}$.
Figure 5: Validation metrics $U^{{\textnormal{ntr}}}$ (hollow markers) and $U^{{\textnormal{hfa}}}$ (colored markers) given by (36)-(37) shown as a function of $\gamma$ for the loss function $\ell^{{\textnormal{loss}}}_{y}(z)=\ell^{{\textnormal{FIVB}}}_{y}(z)$ defined in (30) using parameters which may be: predefined ($\boldsymbol{r}^{{\textnormal{FIVB}}}$), analyticaly calculated ($\tilde{\boldsymbol{r}}\big{(}\boldsymbol{c}^{{\textnormal{FIVB}}}\big{)}$), or, numerically optimized ($\hat{\boldsymbol{r}}$).
The ALO metrics are shown in Fig. 5, where we observe:
  • The ALO metric obtained with optimized numerical scores $\hat{\boldsymbol{r}}$ and with the calculated ones $\tilde{\boldsymbol{r}}(\boldsymbol{c})$ are practically identical, which supports our analysis in Sec. 3.3.1.
  • By comparing the results from Fig. 5 with those shown in Fig. 4 we see that, using the implicit FIVB loss functions, $\ell^{{\textnormal{FIVB}}}_{y}(z)$, and provided the numerical scores $\boldsymbol{r}$ are adequately set (i.e., optimized via (50) or calculated via (46)), a negligible loss of performance is incurred when comparing to the log-score $\ell_{y}(z)$. This justifies the choice made in the FIVB ranking which avoids the exact, but complicated derivative of the loss function shown in (27).
  • The numerical scores $\boldsymbol{r}^{{\textnormal{FIVB}}}$ currently used in the FIVB ranking, lead to the observable performance loss.
The numerical score $\boldsymbol{r}^{{\textnormal{FIVB}}}$ used currently in the ranking is compared, in Fig. 6, to $\hat{\boldsymbol{r}}(\boldsymbol{c}^{{\textnormal{FIVB}}})$ which is obtained by optimization (50) and to $\tilde{\boldsymbol{r}}(\boldsymbol{c}^{{\textnormal{FIVB}}})$ – calculated from (46); since $\boldsymbol{r}^{{\textnormal{FIVB}}}$ and $\tilde{\boldsymbol{r}}(\boldsymbol{c}^{{\textnormal{FIVB}}})$ do not depend on the regularization parameter $\gamma$, they are shown as horizontal lines. Also note that $r_{0}^{{\textnormal{FIVB}}}=\tilde{r}_{0}(\boldsymbol{c}^{{\textnormal{FIVB}}})=\hat{r}_{0}(\boldsymbol{c}^{{\textnormal{FIVB}}})=2$.
The optimized numerical scores $\hat{\boldsymbol{r}}$ change with $\gamma$, but have no incidence on the prediction capacity of the model, as shown already in Fig. 5. In fact, for $\gamma\approx 0.5$, which minimizes the total loss function, we get $\tilde{r}_{1}(\boldsymbol{c}^{{\textnormal{FIVB}}})\approx\hat{r}_{1}(\boldsymbol{c}^{{\textnormal{FIVB}}})$.
Figure 6: Numerical scores: $r_{l}^{{\textnormal{FIVB}}}$ given by the FIVB ranking, $\tilde{r}_{l}(\boldsymbol{c}^{{\textnormal{FIVB}}})$ calculated via (46), and $\hat{r}_{l}(\boldsymbol{c}^{{\textnormal{FIVB}}})$ optimized in (50).
On the other hand, $\hat{r}_{2}(\boldsymbol{c}^{{\textnormal{FIVB}}})$ becomes negative. For example, using the results for $\gamma\approx 0.5$ implies using,
$\hat{r}_{0}$ $=2.0,\quad\hat{r}_{1}\approx 0.9,\quad\hat{r}_{2}\approx-0.1,\quad\hat{r}_{3}\approx 0.1,\quad\hat{r}_{4}\approx-0.9,\quad\hat{r}_{4}=-2,$ (51)
i.e., the numerical score does not decrease monotonically with the outcomes index, $y$.
This may appear surprising and counterintuitive, but only if we interpret the numerical score as related to the order of the outcomes. We should remember that ordinal variables do not have intrinsic numerical values, and numerical scores are parameters that allow us to adjust the form of the implicit loss function $\ell^{\text{FIVB}}_{y}(z)$. With such a perspective, the non-monotonic behavior of $r_{y}$ is allowed.131313 We still want to know, if, for $\hat{\boldsymbol{r}}$ given in (51) (where $\hat{r}_{3}>\hat{r}_{2}$) the implicit loss functions $\ell^{{\textnormal{FIVB}}}_{y}(z;\hat{\boldsymbol{r}})$ remain convex in $z$. For this, it suffices to calculate $\check{r}(z)$ and verify that it is monotonically increasing in $z$. In fact, this is the case for all $\hat{\boldsymbol{r}}$ we obtained. Note that this does not contradict Lemma 1 because the monotonic behavior of $r_{y}$ was a sufficient (but not necessary) condition to ensure the convexity of $\ell^{{\textnormal{FIVB}}}_{y}(z;\boldsymbol{r})$.
We also note that we always obtain more “conventional” (monotonic in $y$) behavior of $\tilde{r}_{y}(\boldsymbol{c})$. Since the optimized scores, $\hat{\boldsymbol{r}}$, and the calculated ones, $\tilde{\boldsymbol{r}}$, do not change the performance, it may be preferable to use the latter.
Immediate conclusion is that, the numerical score $\boldsymbol{r}^{{\textnormal{FIVB}}}$ is inadequately set in the current version of the FIVB ranking. However, it can be easily modified, e.g., using rounded values, $\tilde{r}_{1}=1.0$ and $\tilde{r}_{2}=0.25$.
Thus, it appears that the numerical scores $\boldsymbol{r}^{{\textnormal{FIVB}}}$ were not formally optimized — a conjecture supported by the fact that their origin is not explained in (FIVB, 2024). However, regardless of the origin of $\boldsymbol{r}^{{\textnormal{FIVB}}}$, it is more sound to see the numerical score $\boldsymbol{r}$ as free parameters which allow us to make the implicit loss function $\ell^{{\textnormal{FIVB}}}_{y}(z;\boldsymbol{r})$ “behave” similarly to the optimal logarithmic loss $\ell_{y}(z)$.
Figure 7: Validation metrics $U^{{\textnormal{ntr}}}$ (hollow markers) and $U^{{\textnormal{hfa}}}$ (colored markers) given by (36)-(37) shown as a function of $\gamma$ using the loss function $\ell_{y}(z)$ defined in (14) and different strategies of fixing the weights $\xi_{v}$ which depend on the matches’ categories, including equal weights $\xi_{v}\equiv 1$, $\xi^{{\textnormal{FIVB}}}_{v}$ specified in Table 1, and $\hat{\xi}_{v}$, optimized via (52).
3.4 Weights
The weights $\boldsymbol{\xi}$ are optimized as follows:
$\hat{\boldsymbol{\xi}},\hat{\eta}=\mathop{\mathrm{argmin}}_{\boldsymbol{r},\eta}U(\boldsymbol{r},\eta,\boldsymbol{p}_{\backslash\{\boldsymbol{r},\eta\}}),$ (52)
where we use the thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$ and the log-score function $\ell^{{\textnormal{loss}}}_{y}(z)=\ell_{y}(z)$. The starting point for optimization is $\boldsymbol{\xi}=[1,\ldots,1]$.
The results are shown in Fig. 7, and the conclusion is straightforward: using the weights $\xi_{v}$ specified by the FIVB ranking and given in Table 1, is detrimental to the prediction capacity of the model. The optimization is also practically useless, and, in fact, the optimized weights were quite similar. That is, for $\gamma<0.5$, we obtained $\hat{\xi}_{v}\in(0.9,1.5)$ (not shown here).
To clarify the somewhat intriguing behavior of $U(\boldsymbol{p})$ for the optimized results (marked with stars in Fig. 7), which, for large $\gamma$, is notably better than in the case of $\xi_{v}\equiv 1$, we write the optimization (35) as
$\hat{\boldsymbol{\theta}}_{\backslash{t}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}\Big{[}\sum_{\begin{subarray}{c}\tau=1\\ \tau\neq t\end{subarray}}^{T}\hat{\xi}_{v_{\tau}}(\hat{\gamma})\ell^{{\textnormal{loss}}}_{y_{\tau}}(\boldsymbol{x}_{\tau}^{\mathsf{T}}\boldsymbol{\theta})+\hat{\gamma}\|\boldsymbol{\theta}\|^{2}\Big{]}.$ (53)
where $\hat{\xi}_{v}(\hat{\gamma})$ are the optimal weights obtained in (52) for an optimal $\hat{\gamma}\approx 0.5$. Since the multiplication of the cost function by $\alpha>0$ is irrelevant to the optimization results, using weights $\alpha\hat{\xi}_{v}(\gamma)$ and the regularization parameter $\gamma=\alpha\hat{\gamma}$, will not change $\hat{\boldsymbol{\theta}}_{\backslash{t}}$. In other words, increasing $\gamma$ we can simultaneously increase $\boldsymbol{\xi}$ and maintain performance $U(\boldsymbol{p})$ flat in $\gamma$. The only reason it does not happen is because we impose the constraint $\xi_{0}=1$.
4 Real-time ranking
In previous sections, model evaluation relied on comparing analytically deduced parameters with those obtained through optimization. Our objective now is to directly use the model obtained thanks to analytical insights and to evaluate the performance of the real-time ranking based on the resulting SG algorithm.
We will thus keep the FIVB model defined by the threshold parameters $\boldsymbol{c}^{{\textnormal{FIVB}}}$ and evaluate i) the choice of the numerical score $\boldsymbol{r}$ used in the implicit loss function $\ell^{{\textnormal{FIVB}}}$, and specified in (47) and ii) the values of the weights $\boldsymbol{\xi}$, see Sec. 3.4, Simply put, we do not carry out any explicit optimization of the model parameters when using the real-time ranking but, rather, rely on the parameters obtained from analysis. In this way, we avoid the contentious issue of choosing the model parameters from the data. The only exception is the choice of the HFA parameter which we set as $\eta=0.2$.
To initialize the SG algorithm (10), we use the skills $\theta_{m,0},m=1,\ldots,M$, where $\theta_{m,0}$ is read from the official FIVB ranking of the team $m$ at the time of their first match after 2020.141414Thus, we do not take into account the fact that the FIVB ranking penalizes “inactive” teams, i.e., those that do not play any matches in a given year, and whose ranking is then reduced by 50 points. By initializing the skills with those provided by the official ranking allows us to deal with the practical aspect of switching from one ranking (here, the official one) to another (the one we propose).
We calculate the validation metrics
$\overline{U}$ $=\frac{1}{T}\sum_{t=1}^{T}\ell^{{\textnormal{val}}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}}_{t})$ (54)
$\overline{U}^{{\textnormal{ntr}}}$ $=\frac{1}{T^{{\textnormal{ntr}}}}\sum_{\begin{subarray}{c}t=1\\ h_{t}=0\end{subarray}}^{T}\ell^{{\textnormal{val}}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}}_{t})$ (55)
$\overline{U}^{{\textnormal{hfa}}}$ $=\frac{1}{T^{{\textnormal{hfa}}}}\sum_{\begin{subarray}{c}t=1\\ h_{t}=1\end{subarray}}^{T}\ell^{{\textnormal{val}}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}}_{t}).$ (56)
These metrics are, in essence, equivalent to those in (36)-(37), which we have shown in the figures. The difference is that now skills $\hat{\boldsymbol{\theta}}_{t}$ are estimated using the SG algorithm.
The results obtained are shown in Table 4, where we indicate the loss function used (which determines the gradient used in the SG algorithm) and the parameters of the underlying model.
loss parameters $\mu$ $\overline{U}$ $\overline{U}^{{\textnormal{ntr}}}$ $\overline{U}^{{\textnormal{hfa}}}$ $\overline{\rho}$
A $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0$ 0.01 1.52 1.51 1.53 0.94
0.03 1.49 1.49 1.49 0.89
B $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0.2$ 0.01 1.52 1.51 1.52 0.94
0.03 1.48 1.49 1.47 0.89
C $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\tilde{\boldsymbol{r}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0.2$ 0.01 1.52 1.51 1.53 0.94
0.04 1.47 1.48 1.45 0.88
D $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\xi_{v}\equiv 1,\eta=0.2$ 0.10 1.48 1.49 1.45 0.87
E $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\tilde{\boldsymbol{r}},\xi_{v}\equiv 1,\eta=0.2$ 0.10 1.47 1.48 1.44 0.88
F $\ell$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\xi_{v}\equiv 1,\eta=0.2$ 0.20 1.46 1.48 1.43 0.85
Table 4: The metrics (54)-(56) obtained using the SG algorithm using different loss functions and parameters. We always show the results with the step $\hat{\mu}$ obtained via (57), except in the cases A, B and C, where we also show the results obtained using $\mu=0.01$ which is defined in the FIVB ranking.
For the algorithms based on the FIVB implicit loss function and the weighting with $\boldsymbol{\xi}^{{\textnormal{FIVB}}}$, we evaluate the performance using the nominal adaptation step $\mu=0.01$, and, for each algorithm, we also search for the adaptation step which minimizes the validation metric overall
$\hat{\mu}$ $=\mathop{\mathrm{argmin}}_{\mu}\overline{U}(\mu),$ (57)
where $\overline{U}(\mu)=\overline{U}$ shown in (54).
To indicate how much the new algorithms change the ranking when comparing to the official FIVB ranking, we calculate the average Spearman correlation coefficient
$\overline{\rho}$ $=\frac{1}{T}\sum_{t=1}^{T}\rho(\boldsymbol{\theta}_{t}^{{\textnormal{FIVB}}},\hat{\boldsymbol{\theta}}_{t}),$ (58)
where $\rho(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime})$ is the Spearman correlation between the skill vectors $\boldsymbol{\theta}$ and $\boldsymbol{\theta}^{\prime}$. [^15] If the order implied by the values in $\boldsymbol{\theta}$ is the same as the order implied by $\boldsymbol{\theta}^{\prime}$, we have $\rho(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime})=1$; if, on the other hand, $\boldsymbol{\theta}^{\prime}$ is obtained by taking elements of $\boldsymbol{\theta}$ in reversed order, we have $\rho(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime})=-1$.
Note that even using exactly the same parameters as the FIVB ranking (case A in Table 4), our results are not the same as the official ones because we discarded the forfeited matches; this explains why the Spearman correlation is the largest among the algorithms, but it is not perfect, i.e., $\overline{\rho}<1$.
Without surprise, the best result $\overline{U}=1.46$, is obtained using the true log-score $\ell_{y}(z)$, given by (26) (case F in Table 4) and, on average, this ranking is the least correlated with the FIVB ranking ($\overline{\rho}=0.85$); remember, however, that the implementation of (27) is numerically complex. The second-best result is obtained using i) the numerical scores $\tilde{r}_{y}(\boldsymbol{c}^{{\textnormal{FIVB}}})$, see (46), together with the constant weighting $\xi_{v}\equiv 1$ (case E in Table 4). An improvement in performance from using the HFA $\eta=0.2$ is slight (see cases A and B) but, quantitatively, in line with the improvement observed in home-matches in Fig. 3 (where, after applying the HFA, the value of $U(\boldsymbol{p})$ changes from $1.36$ to $1.34$).
To show an example of how the algorithms (A, E and F) affect the ranking of the top teams, we show the ranking in Table 5, where the three front-runners stay the same, but the position of the remaining teams changes.
case top teams and skills
A ($\mu=0.01$) POL USA JPN BRA ITA ARG RUS
423.8 396.8 345.9 345.0 344.3 317.0 315.7
A ($\mu=0.03$) POL USA JPN ARG SLO GER BRA
517.4 478.2 420.2 374.9 368.1 361.3 360.2
E POL USA JPN ARG ITA SLO GER
468.9 462.2 409.8 381.0 375.3 366.4 356.0
F POL USA JPN GER ARG SLO ITA
528.9 524.3 475.3 436.9 430.6 416.3 401.4
Table 5: Ranking of the top teams in the last day of 2023 for different algorithm with parameters specified in Table 4.
Regarding the choice of the adaptation step, we observe that
  • The adaptation step $\mu=0.01$ used by the FIVB ranking is too small and, by increasing it three- or four-fold, performance improves. This can be explained using the interpretation of the SG algorithm as a simplified Kalman filter proposed by Szczecinski and Tihon (2023, Sec. 3.3), where the adaptation step in the SG algorithm has a meaning of the posterior variance of the skills. In simple terms, FIVB is over-optimistic about the uncertainty (variance) in the estimation of the skills.
  • Since the weighting may also be interpreted as a variable step size, by removing it, i.e., using $\xi_{v}\equiv 1$, we have to explicitly increase the step size; this explains the large value of $\mu$ for each configuration in which we use $\xi_{v}\equiv 1$.
5 Conclusions
In this work, the online ranking algorithm used by the FIVB is presented in the statistical learning framework. To our best knowledge, the FIVB ranking is the first to adopt an explicit probabilistic model (here, the Cumulative Link (CL)model) of the multi-level ordinal outcomes, and, from the statistical perspective, this is a step in the right direction. On the other hand, the algorithms adopt simplifications that we demonstrate to be suboptimal, which is the “misstep” in the title. However, we show how these simplifications may be easily corrected using well-defined formulas to calculate the numerical scores, see (46). The impact of these changes on the on-line ranking can be seen in Table 4 and Table 5.
To analyze the algorithm, we use two approaches: i) the analytical, where the approximations and simplifications allow us to draw conclusions about the properties of the model, as well as to optimize its parameters, and ii) the numerical, where we explicitly optimize the parameters of the model from the outcomes of the international volleyball matches used in the FIVB ranking.
The analytical approach is easily reproducible, while the numerical optimization which relies on the cross-validation strategy allows us to validate the insights obtained analytically. This led to the following understanding of the current FIVB ranking algorithm:
  • The FIVB algorithm should be seen as the approximate ML inference of the skills from the ordinal match outcomes. The approximations are due to the use of the SG to solve the optimization problem, and, more importantly, due to the use of the loss functions, which are proxies for the log-likelihood of the ML approach. We explain the rationale for using such proxy loss functions.
  • Although the form of loss functions is not explicitly mentioned in the description of the FIVB algorithm, they can be inferred from the equations, and we show how they depend on the numerical scores that are attributed to the ordinal match outcomes in the FIVB algorithm. This is interesting because in this way we explain the meaning of the numerical scores, as, from the modeling perspective, the ordinal variables do not have numerical values.
Regarding the model underlying the current FIVB ranking algorithm and the algorithm itself, we studied:
  • CL model thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$, see (18), which define the probabilistic ordinal model of the data. They fit relatively well the data we analyzed (the FIVB matches from 2021-2023), see Sec. 3.2.
  • Numerical scores $\boldsymbol{r}^{{\textnormal{FIVB}}}$, see Table 2, which define the algorithm. These are shown, both analytically and numerically, to be inadequately set, see Sec. 3.3.
  • Importance weights $\boldsymbol{\xi}$, see Table 1, which change the contribution of the outcome of the match depending on the match type. These are shown to be irrelevant from an statistical point of view, see Sec. 3.4.
  • Home-field advantage (HFA) which deals with the matches played on the home venues by artificially boosting the skills of the home-team. We found that the HFA is a relevant parameter that improves the prediction performance. The improvements are relatively small, which may explain why the current FIVB ranking does not use the HFA. On the other hand, there is no cost related to its application.
Recommendations
In summary, by keeping the structure of the current FIVB ranking and by exploring the optimality of the above model/algorithm choices, we came up with new parameters that can improve the performance of the algorithm. And, since FIVB explicitly says that its algorithm may be updated, if this is to happen, our recommendations, in order of importance, are the following:
  1. 1.
    Change the numerical score $\boldsymbol{r}^{{\textnormal{FIVB}}}$ to be similar to those suggested in (47)-(49).
  2. 2.
    Introduce the HFA to the algorithm. This is a simple and low-cost modification, yielding an improvement in the prediction of the matches played on the home-venue.
  3. 3.
    Remove the weighting of the matches with $\xi^{{\textnormal{FIVB}}}_{v}$. Or, if its use is motivated by some extra-statistical (e.g., entertainment) reasons, decrease the differences between the possible values of $\xi_{v}^{{\textnormal{FIVB}}}$.
Further work
While in this work, we focus on the FIVB ranking, the evaluation methodology we propose can be used more broadly, to analyze the ranking algorithms. In particular, the “reverse engineering” approach we used to reveal the form of the (implicit) objective loss function (see Sect. 2.5) is particularly useful. It can be applied to analyze the sub-optimality of the ranking algorithms which, in practice, may be defined without an explicit probabilistic model. In that sense, our evaluation methodology is more general than reverse engineering, which was used in Szczecinski and Roatis (2022) to unveil the model underlying the FIFA ranking.
Beyond this general recommendation and focusing specifically on improving the FIVB ranking, the following venues can be explored:
  • Considering a time-variant model for the skills in the design of the algorithm, e.g., using ideas already shown before in (Fahrmeir, 1992), (Glickman, 1993), (Knorr-Held, 2000), (Szczecinski and Tihon, 2023). It may require particular attention, as the current version of the FIVB ranking algorithm is not a straightforward implementation of the ML ranking.
  • Analyzing alternative ordinal models McCullagh (1980), taking into account the simplicity of the algorithm they produce, including the use of different CDF in the model (16) (Tutz, 2012, Ch. 9.1.3)(Agresti, 2013, Ch. 8.3), or the Adjacent Categories (AC) models (Tutz, 2012, Ch. 9.4.5) Szczecinski (2022).
Appendix A Notation in FIVB ranking
We show in Table 6, the relationship between our notation and the one used in the FIVB ranking description (FIVB, 2024), where the following abbreviations are used
  • WR: World ranking (here, a FIVB ranking)
  • WRS: World ranking score (here, we call it skills $\boldsymbol{\theta}_{m,t}$)
  • SSV: Set score variant (we call it numerical score $r_{y}$)
  • EMR: Expected match result (here, the expected score $\check{r}(z)$, see (22))
  • MWF: Match weighting factor (here, it corresponds to $10\xi_{v_{t}}$)
[…]
Our notation FIVB notation
$\theta_{m,t},\theta_{n,t}$ WRS1, WRS2
$c_{0}^{{\textnormal{FIVB}}},\ldots,c_{4}^{{\textnormal{FIVB}}}$ C1, …, C5
[…]
$s$ 1000/8=125
$\mathsf{P}_{0}(z_{t}),\ldots,\mathsf{P}_{5}(z_{t})$ P1, …, P5
$\Phi(z)$ $\sim N(0,1)(z)$
$r_{y_{t}}^{{\textnormal{FIVB}}}$ SSV
$\check{r}(z_{t}/s)$ EMR
$10\xi_{v_{t}}$ MWF
$-g^{{\textnormal{FIVB}}}(z_{t}/s)=r_{y_{t}}^{{\textnormal{FIVB}}}-\check{r}(z_{t}/s)$ WR value
$\theta_{m,t+1}-\theta_{m,t}=-10\xi_{v_{t}}g^{{\textnormal{FIVB}}}(z_{t}/s)$ WR points = WR values * MWF /8
Table 6: Equivalence of this work’s notation and the one used in the description of the FIVB ranking.
With this notation, the update formula is given by
[…]
and corresponds to (20) which, focusing on the update of the skills of the home team $m$, may be written as
$\theta_{m,t+1}$ $=\theta_{m,t}-\mu s\xi_{v_{t}}g^{{\textnormal{FIVB}}}_{y_{t}}(z_{t}/s).$ (60)
Appendix B Optimization of the cross-validation metric
The simplest optimization of the cross-validation metric $U(\boldsymbol{p})$ may be done via the steepest descent
$\hat{\boldsymbol{p}}\leftarrow\hat{\boldsymbol{p}}-\kappa\nabla_{\boldsymbol{p}}U(\boldsymbol{p}),$ (61)
where, $\kappa$ is the step-size, and to calculate the gradient $\nabla_{\boldsymbol{p}}U(\boldsymbol{p})$, we have to calculate derivatives of $U(\boldsymbol{p})$ with respect to a parameter $q\in\boldsymbol{p}$. This can be done as follows:
$\frac{\partial}{\partial q}U(\boldsymbol{p})$ $=\frac{1}{T}\sum_{t=1}^{T}\frac{\partial}{\partial q}\ell^{{\textnormal{val}}}_{y_{t}}(\hat{z}_{t,\backslash{t}},\boldsymbol{p})$ (62)
$\frac{\partial}{\partial q}\ell^{{\textnormal{val}}}_{y_{t}}(\hat{z}_{t,\backslash{t}})$ $=\frac{\partial\hat{z}_{t,\backslash{t}}}{\partial q}\dot{\ell}^{{\textnormal{val}}}(\hat{z}_{t,\backslash{t}},\boldsymbol{p})+\frac{\partial}{\partial q}\ell^{{\textnormal{val}}}_{y_{t}}(\hat{z}_{t,\backslash{t}},\boldsymbol{p})$ (63)
$\frac{\partial\hat{z}_{t,\backslash{t}}}{\partial q}$ $=\frac{\partial\hat{z}_{t}}{\partial q}+\frac{\partial}{\partial q}\left[\frac{\xi_{v_{t}}\dot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})a_{t}}{1-\xi_{v_{t}}\ddot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})a_{t}}\right].$ (64)
In (64), we will need
$\frac{\partial\hat{z}_{t}}{\partial q}$ $=\boldsymbol{x}^{\mathsf{T}}_{t}\frac{\partial\hat{\boldsymbol{\theta}}}{\partial q}$ (65)
$\frac{\partial a_{t}}{\partial q}$ $=\boldsymbol{x}^{\mathsf{T}}_{t}\frac{\partial\hat{{\textnormal{{H}}}}^{-1}}{\partial q}\boldsymbol{x}_{t}=-\boldsymbol{x}^{\mathsf{T}}_{t}\hat{{\textnormal{{H}}}}^{-1}\frac{\partial\hat{{\textnormal{{H}}}}}{\partial q}\hat{{\textnormal{{H}}}}^{-1}\boldsymbol{x}_{t}$ (66)
$\frac{\partial\hat{\textnormal{{H}}}}{\partial q}$ $=\sum_{t=1}^{T}\boldsymbol{x}_{t}\frac{\partial}{\partial q}\big{[}\xi_{v_{t}}\ddot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})\big{]}\boldsymbol{x}^{\mathsf{T}}_{t}+\mathds{I}\big{[}{q=\gamma}\big{]}{\textnormal{{I}}}$ (67)
$\frac{\partial}{\partial q}\dot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})$ $=\frac{\partial\hat{z}_{t}}{\partial q}\ddot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})+\frac{\partial}{\partial q}\dot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})$ (68)
$\frac{\partial}{\partial q}\ddot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})$ $=\frac{\partial\hat{z}_{t}}{\partial q}\dddot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})+\frac{\partial}{\partial q}\ddot{\ell}_{y_{t}}(\hat{z}_{t},\boldsymbol{p})$ (69)
where, in (66) we used (Petersen and Pedersen, 2012, Eq. (40)), and ˙˙˙$\dddot{\ell}_{y}(z,\boldsymbol{p})=\frac{\partial^{3}}{\partial z^{3}}\ell_{y}(z,\boldsymbol{p})$.
From implicit function theorem, see (Lorraine et al., 2019, Th. 1), using $\hat{\boldsymbol{\theta}}\equiv\hat{\boldsymbol{\theta}}(\boldsymbol{p})$
$\boldsymbol{0}$ $=\frac{\partial}{\partial q}\Big{[}\nabla_{\boldsymbol{\theta}}J(\hat{\boldsymbol{\theta}}(\boldsymbol{p}),\boldsymbol{p})\Big{]}$ (70)
$\boldsymbol{0}$ $=\nabla^{2}_{\boldsymbol{\theta}}J(\hat{\boldsymbol{\theta}}(\boldsymbol{p}),\boldsymbol{p})\frac{\partial\hat{\boldsymbol{\theta}}(p)}{\partial q}+\frac{\partial}{\partial q}\nabla_{\boldsymbol{\theta}}J(\hat{\boldsymbol{\theta}},\boldsymbol{p})$ (71)
$\frac{\partial\hat{\boldsymbol{\theta}}}{\partial q}$ $=-\hat{{\textnormal{{H}}}}^{-1}\frac{\partial}{\partial q}\nabla_{\boldsymbol{\theta}}J(\hat{\boldsymbol{\theta}},\boldsymbol{p})$ (72)
$\frac{\partial}{\partial q}\nabla_{\boldsymbol{\theta}}J(\hat{\boldsymbol{\theta}},\boldsymbol{p})$ $=\sum_{t=1}^{T}\frac{\partial}{\partial q}\big{[}\xi_{v_{t}}\dot{\ell}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}},\boldsymbol{p})\big{]}\boldsymbol{x}_{t}+\mathds{I}\big{[}{\gamma=q}\big{]}\hat{\boldsymbol{\theta}}.$ (73)
By plugging $\hat{z}_{t,\backslash{t}}(\boldsymbol{p})$ into (33) we obtain the function $U(\boldsymbol{p})$ which depends on $\boldsymbol{p}$ and we can calculate the gradient $\nabla_{\boldsymbol{p}}U(\boldsymbol{p})$.
Similarly, we can use the Newton method
[…]
where, to calculate the Hessian, $\nabla^{2}_{\boldsymbol{p}}U(\boldsymbol{p})$ we need second order derivatives.
However, the expressions for the gradient and, especially, the Hessian, quickly become cumbersome; see (Burn, 2020). Thus, instead of explicit differentiation, we use the automatic differentiation available in JAX and JAXopt python-compliant packages (Bradbury et al., 2018), (Blondel et al., 2021) with a particularly interesting feature which automatically finds the implicit differentiation required to find the derivative of $\hat{\boldsymbol{\theta}}$ with respect to hyperparameters in $\boldsymbol{p}$ as specified by (72).
We do not show more details to not overcomplicate the presentation, especially that they are not really required because the numerical optimization is used to confirm the observation we made using the analytical insight. In fact, the performance of the online algorithm shown in Sec. 4 uses the parameters (shown in Table 4) which are explicitly defined prior to the application of the algorithms.